### v1

# Integrated modelling of equilibrium and transport in axisymmetric magnetic mirror fusion devices

Abstract
This paper presents the Hammir tandem mirror design based on Realta Fusion’s first-of-a-kind model for axisymmetric magnetic mirror fusion performance. This model uses an integrated end plug simulation model including, heating, equilibrium, and transport combined with a new formulation of the plasma operation contours (POPCONs) technique for the tandem mirror central cell. Using this model, it is shown that an end plug utilizing high temperature superconducting magnets and modern neutral beams enables a classical tandem mirror pilot plant producing a fusion gain Q > 5. The approach here represents an important advance in tandem mirror design. The high fidelity end plug model enables calculations of heating and transport in the highly non-Maxwellian end plug to be made more accurately and the central cell POPCON technique allows consideration of a wide range of parameters in the relatively simple near-Maxwellian central cell, facilitating the selection of more optimal central cell plasmas. These advances make it possible to find more conservative classical tandem mirror fusion pilot plant operating points with lower $\beta$, temperatures, neutral beam energies, and end plug performance than designs in the literature. Despite being more conservative, it is shown that these operating points can still form the basis of a viable fusion pilot plant.
1 The compact mirror approach to fusion
[…]
The new magnetic mirror development program being undertaken by Realta Fusion takes advantage of a number of key enabling technology and physics breakthroughs that can greatly enhance the performance of the magnetic mirror to revive the tandem mirror fusion plant concept: high-energy negative-ion neutral beams (Kashiwagi et al., 2021; Serianni et al., 2023), modern high-frequency gyrotrons (Rzesnicki et al., 2022; ARPA-E, 2023), HTS magnets (Michael et al., 2017; Whyte, 2019), and magnetohydrodynamic (MHD) stabilization of axisymmetric mirrors (Beklemishev et al., 2010; Ryutov et al., 2011; Endrizzi et al., 2023). These advancements allow less complex magnetic mirrors to be constructed with planar HTS coils that reach substantially higher fields and mirror ratios than previous magnetic mirror experiments and can dramatically improve tandem mirror performance (Fowler et al., 2017). Presently, Realta is sponsoring research at the University of Wisconsin-Madison on the Wisconsin High-Temperature Superconducting Axisymmetric Magnetic Mirror (WHAM) experiment, an axisymmetric mirror using two custom $17~{}\mathrm{T}$ HTS magnetic field coils built by Commonwealth Fusion Systems. A comprehensive description of the WHAM experiment is available in Endrizzi et al. (2023). The next step device, Anvil, will be a large simple mirror constructed by Realta Fusion to serve as an end plug physics demonstrator and fusion technology test platform. Anvil will be a similar class of device to the break-even axisymmetric mirror (BEAM) described in Forest et al. (2024). The Realta Fusion tandem mirror pilot plant Hammir is the final step on Realta Fusion’s technology development roadmap and will be the focus of this paper. Hammir will be a smaller, more conservative, axisymmetric tandem mirror than those identified in the Fowler et al. (2017) analysis, which considered only axisymmetric tandem mirrors with $\beta\sim 1$, very high temperatures $T_{i}=T_{e}=150~{}\mathrm{keV}$, high energy $1~{}\mathrm{MeV}$ neutral beams, and large fusion power outputs $P_{fus}>1~{}\mathrm{GW}$. It will be shown in this work that smaller classical tandem mirrors with significantly lower $\beta$, temperature, beam energies, and fusion power are still competitive with other fusion concepts and compatible with the requirements for a commercial fusion plant identified by the National Academies of Sciences, Engineering, and Medicine (NASEM) “Bringing Fusion to the U.S. Grid” report (Hawryluk et al., 2021), specifically:
  • Electrical gain: $Q_{e}=P_{\mathrm{ele,out}}/P_{\mathrm{ele,in}}>1$
  • Continuous net electricity $P_{\mathrm{ele,out}}>50~{}\mathrm{MWe}$ for at least 3 hours (or the process heat equivalent in the case of a industrial process heat use case)
To simplify tandem mirror modelling, this work takes advantage of the properties of the classical tandem mirror to split up the modelling problem. The central cell and the end plug are handled using separate models that are then coupled together. The complicated non-Maxwellian end plug which undergoes intense heating is analyzed using a high fidelity integrated simulation model, whereas, the simpler near-Maxwellian central cell is extrapolated based on the end plug parameters using a novel version of the Plasma OPeration CONtour (POPCON) technique (Cordey et al., 1981; Houlberg et al., 1982). This approach allows larger parameter spaces to be considered and breaks the tandem mirror design problem down to a few key engineering and physics parameters which are important to design such as the plasma radius at the high field mirror coil throat $a_{m}$, the field at the mirror throat $B_{m}$, and the end plug density $n_{p}$. These are key steps to developing effective pilot plant designs, and successful tokamak designs also have used multifidelity models and identified analogous design parameters such as the minor radius $a$, major radius $R$, on axis field $B_{0}$, and plasma current $I_{p}$ (Sorbom et al., 2015; Creely et al., 2020; Buttery et al., 2021; Frank et al., 2022b; The MANTA Collaboration et al., 2024).
The rest of this paper is structured as follows: Section 2 describes the new RealTwintm integrated model for a Hammir end plug. Section 3 determines what end plug parameters are required for a viable Hammir fusion power plant using a central cell POPCON. Section 4 identifies an optimal operating point for the Hammir end plug which satisfies the end plug requirements identified by the POPCONs in Section 3 under realistic physics and engineering constraints using the RealTwin model. Finally, Section 5 summarizes the work, identifying key capability gaps discovered during the Hammir design process and plans to address them in the future.
2 The RealTwin simulation model
2.1 Fokker-Planck transport and plasma heating using CQL3D-m and Freya
Mirror transport during high performance operation where MHD flute modes and kinetic modes such as the drift-cyclotron loss-cone (DCLC) instability are not present is dominated by axial transport from collisions and neutral dynamics. To simulate transport under these conditions, a Fokker-Planck equation solver including plasma heating and fueling models is employed. This work used the CQL3D-m Fokker-Planck code and its internal neutral beam deposition Monte-Carlo solver Freya.
The CQL3D-m code (Harvey et al., 2016) used here is a modified version of the well known CQL3D Fokker-Planck code (Harvey & McCoy, 1992) and older versions of CQL3D-m were used for the design of WHAM and the BEAM devices (Endrizzi et al., 2023; Forest et al., 2024). Much like CQL3D, CQL3D-m solves the bounce-averaged Fokker-Planck equations for a coupled set of plasma species $\alpha,\beta,...$. The Fokker-Planck equation for plasma species $\alpha$ is:
$\frac{df_{\alpha 0}}{dt}=C(f_{\alpha 0},f_{\beta 0},...)+Q(f_{\alpha 0},\boldsymbol{E})+R(f_{\alpha 0})+S_{+}(f_{\alpha 0})-S_{-}(f_{\alpha 0}),$ (1)
where $f_{\alpha 0}$ is the bounce-averaged distribution function of species $\alpha$ at the midplane, where $z=0$ (values taken at the midplane are denoted with subscript 00), $f_{\beta 0}$ is the distribution function of species $\beta$ etc., $C(f_{\alpha 0},f_{\beta 0},...)$ is the bounce-averaged nonlinear form of the Rosenbluth collision operator (Rosenbluth et al., 1957) considering collisions between all species in the simulation using their fully numerical distribution functions, $Q(f_{\alpha 0},\boldsymbol{E}_{RF})$ is the bounce-averaged quasilinear diffusion coefficient associated with radiofrequency (RF) electric field $\boldsymbol{E}$ arising from either RF heating or an instability such as the DCLC, $R(f_{\alpha 0})$ is the radial diffusion (not considered in this work as losses from radial diffusion are ideally much smaller than axial losses in the end plug), and $S_{+}$ and $S_{-}$ are particle source and sink terms respectively arising from mirror losses, neutral fueling/charge-exchange, etc.
CQL3D-m discretizes (1) in two velocity-space dimensions, $u=\gamma v$ the momentum-per-rest-mass (where $v$ is particle velocity and $\gamma$ is the relativistic factor) and $\vartheta$ the velocity-space pitch angle, as well as one spatial dimension $\psi$ the square-root poloidal flux (a radial-like coordinate in the magnetic mirror), and uses the implicit time advance solver described in Kerbel & McCoy (1985). The coefficients $Q$, $C$, $R$, $S_{\pm}$ in (1) have been bounce-averaged in magnetic mirror geometry assuming zero-orbit-width. The bounce-average of some coefficient $A$ is defined:
$\langle\langle A\rangle\rangle=\frac{1}{\tau_{b}}\oint\frac{d\ell_{b}}{\left|v_{\parallel}\right|}A\Bigg{|}_{\mu_{m},\vartheta_{0},\psi}.$ (2)
The mirror loss boundaries in CQL3D-m are calculated by the following method. Particles starting with a certain $\vartheta_{0}$ and $u_{0}$ at the midplane can be trapped by either the magnetic mirror force or the mirror’s ambipolar potential. These two trapping mechanisms have distinct loss boundaries that can be computed by the following methods. Potential trapping occurs when for species $\alpha$, the quantity $v_{p}^{2}(z)$ becomes less than zero at some point in $z$ along the flux surface. The definition of this quantity is:
$v_{p}^{2}(z)=2\frac{q_{\alpha}}{m_{\alpha}}(\phi_{p0}-\phi_{p}(z)),$ (3)
where $\phi_{p}(z)$ is the Pastukhov potential (Pastukhov, 1974; Cohen et al., 1978). This can be used to define a potential trapping boundary where:
$v_{\phi_{p}}(z)=\sqrt{-v_{p}(z)^{2}}.$ (4)
For a species where $v_{p}^{2}>0$ along the entire flux surface in $z$, no potential trapping takes place. Similarly, the mirror trapping boundary for species $\alpha$ with $v_{0}$ can be computed with equations:
$v_{\parallel,\mathrm{mirr}}$ $=v_{0}\sin{\vartheta_{\mathrm{mirr}}}$ (5a)
$\vartheta_{\mathrm{mirr}}$ $=\arcsin{\left(\sqrt{\frac{B(z)}{B_{0}}\frac{v_{0}^{2}}{v_{0}^{2}+v_{p}(z)^{2}}}\right)_{\mathrm{max},z}},$ (5b)
where the quantity in parentheses is evaluated at its maximum value in $z$ along a poloidal flux surface. If $\phi_{p}(z)$ is taken to be zero everywhere, this becomes the standard expression for the mirror loss cone. Components of the distribution function $f$ which have velocities greater than the outermost loss boundary, potential or mirror, in a simulation are removed after the time advance. The loss boundaries are then recomputed for the next timestep based on an updated $\phi_{p}(z)$ profile calculated with the method below.
CQL3D-m accounts for ambipolarity preservation by directly calculating the Pastukhov potential response (Pastukhov, 1974; Cohen et al., 1978). This is done through an iterative relaxation solver which determines the potential required to meet the quasineutrality condition $n_{e}=\sum_{si}Z_{si}n_{si}$. The Pastukhov potential at a given $z$ location along a flux surface is computed by iterating two different calculation methods. The first method is an iterative solver which takes:
$\phi_{p}^{(n)}(z)=\phi_{p}^{(n-1)}(z)+f_{\mathrm{rlx}}T_{e0}\frac{\sum_{si}Z_{si}n_{si}(z)-n_{e}(z)}{n_{e}(z)+\sum_{si}Z_{si}n_{si}(z)},$ (6)
where $\phi_{p}^{(n)}(z)$ is the updated Pastukhov potential at the new timestep, $\phi_{p}^{(n-1)}$ is the Pastukhov potential at the previous timestep, $f_{\mathrm{rlx}}\approx 0.5-1.0$ is a relaxation factor added to improve numerical stability, and $n_{s}$ refers to the local value of density along $z$ for species $s$. If only (6) is used to calculate $\phi_{p}(z)$, the code will eventually become unstable. To counteract this, the potential is occasionally relaxed with a Boltzmann like response:
$\phi_{p}(z)=T_{e0}\ln{\left(\frac{n_{e}(z)+\sum_{si}Z_{si}n_{si}(z)}{n_{e0}+\sum_{si}Z_{si}n_{si0}}\right)},$ (7)
which has been found empirically to ensure stability over many time steps. The generation of $\phi_{p}(z)$ leads to off-midplane trapping due to phenomena such as sloshing ions obtained by skew injected neutral beams (Kesner, 1980) or sloshing electrons from directed electron cyclotron heating (ECH) (Baldwin & Logan, 1979). Such trapping is not included in a typical bounce-averaged formulation of the Fokker-Planck equation used by CQL3D-m. To account for these trapped populations, a Maxwellian component is added to the local off-midplane distribution function with a temperature corresponding to the local potential well depth if an insufficient density can be sourced from the remapped midplane distribution function.
The neutral beam solver Freya is unchanged from the typical version implemented in CQL3D and described in Harvey & McCoy (1992); Rosenberg et al. (2004). Neutral particles are injected using realistic neutral beam geometry and tracked until they are ionized in the plasma. The particle source list from Freya is used to calculate a bounce-averaged particle source of ions and electrons with the neutral beam velocity. A bounce-averaged particle sink corresponding to NBI charge exchange is also formulated. Similarly, the computation of the quasilinear diffusion coefficients from ray data in CQL3D-m works in effectively the same way as it did in CQL3D (Harvey & McCoy, 1992) with a few adjustments to the bounce-averaging procedure to account for mirror geometry.
Figure 1: An example of WHAM plasma profiles and distribution functions calculated with CQL3D-m. On the left is a plot from CQL3D-m with logarithmic contours of the ion distribution in arbitrary units for the innermost $\sqrt{\psi_{n}}=0.01$ normalized square root poloidal flux surface with the loss boundary shown in red. On the right are the electron (solid red) and ion (dashed red) density profiles as well as the Pastukhov potential (blue) versus axial distance along the mirror $z$.
An example of a CQL3D-m simulation of WHAM showing a sloshing ions distribution is shown in Figure 1. This provides an example of many of the new features that are present in CQL3D-m. The plasma is fueled by a $45^{\circ}$ NBI producing a sloshing-ions distribution function with strong density peaking about the off-midplane fast ion turning points at $B/B_{0}\approx 2$. The peaking causes a corresponding off-axis potential peak to enforce quasineutrality as the electrons are still effectively Maxwellian which in turn generates a midplane centered potential well which traps lower-energy warm-ions near the midplane.
2.2 MHD equilibrium using Pleiades
In order to calculate the equilibrium magnetic fields of the high beta plasmas under consideration in this work, a custom fork of the Green’s function based magnetic field solver code Pleiades introduced in Peterson et al. (2019) was developed. Pleiades was used in a number of previous works on mirrors (Egedal et al., 2022; Endrizzi et al., 2023; Forest et al., 2024), but the anisotropic equilibrium solver and self-consistent coupling found in this work is represents a significant evolution of the code (these works included diamagnetic corrections to the magnetic equilibrium due to the plasma pressure but used a different method to solve for the corrections, and the corrections were not made self-consistently). This version of Pleiades solves for the MHD momentum balance in flux coordinates and outputs standard format eqdsk files containing the MHD equilibrium. The MHD momentum balance in equilibrium is defined:
[…]
where $\boldsymbol{J}$ is the current density, $\boldsymbol{B}$ is the magnetic field, and $\underline{\underline{\boldsymbol{P}}}$ is the anisotropic pressure tensor. To solve (8) using this technique, the divergence of the anisotropic pressure tensor is rewritten in component form:
[…]
where $\hat{\boldsymbol{b}}=\boldsymbol{B}/|B|$. Taking the cross product of $\boldsymbol{B}$ with both sides of (9) after some vector algebra yields:
[…]
To solve the above equation in Pleiades, the correct choice of flux coordinates must be made such that derivatives of $p$ and $B$ are smooth and can be taken accurately. For an axisymmetric mirror, the most convenient choice of orthogonal coordinates for these purposes is the canonical Clebsch coordinate system ($\psi$, $\phi$, $s$) (D’haeseleer et al., 1991). In this coordinate system, the poloidal flux $\psi$ is the radial-like coordinate, the poloidal angle $\phi$ is the symmetry direction, and the distance along the field line $s$ is directed in the direction of the magnetic field $\hat{\boldsymbol{b}}$. Applying these coordinates to the problem dramatically simplifies (10) providing an equation for the diamagnetic current response $J_{\mathrm{dia}}\equiv J_{\perp}\equiv J_{\phi}$ given the spatial profiles of kinetic $p$ and $B$:
$J_{\mathrm{dia}}=\left[1+\frac{\mu_{0}(p_{\perp}+p_{\parallel})}{B^{2}}\right]^{-1}\left(\frac{1}{B}\frac{\partial p_{\perp}}{\partial\psi}+\frac{p_{\parallel}-p_{\perp}}{B^{2}}\frac{\partial B}{\partial\psi}\right).$ (11)
With (11), Pleiades can be used to calculate the magnetic equilibrium by the following method: first, Pleiades is run without the plasma included; the vacuum fields and poloidal flux, $\boldsymbol{B}_{\mathrm{vac}}(R,Z)$ and $\psi_{\mathrm{vac}}(R,Z)$, produced by the current sources magnetic field coils are solved for using the Green’s functions corresponding to the magnetic field coils’ currents. After the vacuum fields are determined, the anisotropic pressure profiles given on the $(\psi,s)$ grid from CQL3D-m are mapped onto their corresponding $(\psi,s)$ location in the initial Pleiades solution. The value of $J_{\mathrm{dia}}$ is calculated using (11) at each location and the grid centered area $\Delta A(\psi,s)$ is multiplied by the local value of $J_{\mathrm{dia}}(\psi,s)$ to provide a current source $I_{\mathrm{dia}}(\psi,s)$. The Green’s functions associated with the $I_{\mathrm{dia}}(\psi,s)$ current sources are calculated and used along with previously calculated Green’s functions associated with the magnetic coils to compute updated profiles of $\psi$ and $\boldsymbol{B}$ which account for diamagnetism. The anisotropic pressure and its derivatives are then mapped to the updated $\psi$ flux surface locations and the process is repeated until the change in $\psi$ over an iteration drops below some prescribed value $\epsilon=\Delta\psi/\psi$. For $\epsilon=10^{-6}$, this typically takes $\sim 10$ iterations. This flux advance and current redeposition scheme is very similar to the magnetic equilibrium solution scheme used in non-inductive current drive calculations with the tokamak code ACCOME (Devoto et al., 1992). As the edge pressure profile is not solved for in the current version of the RealTwin, an exponential pressure drop at the plasma edge with an $e$-folding length equal to the ion Lamor radius $\rho_{i}$ is used. If the value of $\epsilon$ is made small enough and a sufficiently refined ($\psi$, $s$) grid is used in the computation, the solution to the magnetic equilibrium obtained with this method will converge to the free-boundary pressure balance solution. While the Green’s function approach used by Pleiades is simpler to implement than a free-boundary Grad-Shafranov solver, it is more computationally expensive as Green’s functions corresponding to the diamagnetic current response must be recalculated each time $\psi$ is advanced within the code. However, as each Green’s function computation is independent, the diamagnetic currents and their associated Green’s functions can be computed in parallel, so a refined $(\psi,s)$ grid can be used without incurring long computational times.
Figure 2: An example of a magnetic equilibrium for a high $\beta$ plasma in WHAM calculated with Pleiades. On the left are the flux contours for the vacuum fields $\psi_{\mathrm{vac}}$ in blue and the flux contours after diamagnetic evolution $\psi_{\mathrm{new}}$ in red. On the right are the kinetic pressure profiles and magnetic pressure profiles as well as the paraxial equilibrium condition in green which in paraxial equilibrium is equal to $B_{\mathrm{vac}}^{2}/2\mu_{0}$.
The magnetic equilibrium solver in Pleiades has been quite successful in practice and is robust to high values of $\beta\sim 0.8$. An example of a magnetic equilibrium in WHAM solved using Pleiades appears in Figure 2 and shows the anticipated outward radial expansion of the flux surfaces consistent with a diamagnetic reduction in central magnetic field equal to $B_{0}=B_{\mathrm{vac}}\sqrt{1-\beta}$. The paraxial equilibrium condition from Ryutov et al. (2011), also shown in this figure, is likewise well satisfied (though not perfectly as it is only approximate). The Pleiades simulations used here can be extended to magnetic equilibria with strong plasma flows, i.e. when Mach number $M=\sqrt{\rho v^{2}/(\gamma p)}\gtrsim 1$, where $v$ is the flow velocity, $\gamma$ is the adiabatic constant, and $\rho$ is the mass density of the plasma, by modifying (10) as follows (Endrizzi et al., 2023):
[…]
2.3 Simulation coupling and future components
Figure 3: An example of an IPS driven workflow for calculating simple mirror plasma performance. After each code in the iteration is run the plasma state is updated and used to setup the next simulation.
In order to couple the codes described here together in a self-consistent fashion, a modified version of the Integrated Plasma Simulator (IPS) (Elwasif et al., 2010) was created with wrappers specifically tailored to mirror relevant codes and an HDF5 based mirror plasma state file. The IPS is designed to enable many codes to be run in a single workflow using a driver program that coordinates the execution of different codes, manages their resource usage, and supports advanced features such as concurrent and nested execution of simulations. When each code runs, it interfaces with the IPS through the centralized plasma state file allowing different codes and models to be swapped in and out of the workflow in a modular fashion. This has made the IPS the tool of choice for many public fusion research initiatives including ARPA-E projects and SciDACs (Bonoli et al., 2018; Badalassi et al., 2023; Collins et al., 2023) and has driven large scale coupled simulations using CQL3D previously (Frank et al., 2022a, 2024). Usage of the IPS enables many different physics codes to be coupled together in an efficient manner enabling integrated simulation of mirror plasmas in which the right code can be used for the right job. It also facilitates the level of repeatability and automation needed to do optimizations like the one presented here.
An IPS simulation is initialized with defeatured engineering information about the mirror system under consideration. From that information, the IPS generates a plasma state file which contains the heating system parameters, the plasma parameters, and the magnetic coil parameters as well as the magnetic equilibrium. Transport simulations, like the ones reported on here, are initialized assuming a uniform breakdown plasma (the parameters used for the initial plasma have not been found to notably affect the simulations’ results at steady-state with the exception of the situation in which the initial density is too low to provide significant NBI absorption). The first simulation executed is generally Pleiades to provide a calculation of the vacuum magnetic fields for a given set of coil parameters. Heating sources from RF and/or neutral beams (if an external neutral beam code like FIDASIM is used rather than the Freya solver in CQL3D) are then computed, and a transport simulation with CQL3D-m is run for $\tau\lesssim\tau_{p}\sim 1~{}\mathrm{s}$. The updated plasma profiles computed by CQL3D-m are used to rerun Pleiades to provide an updated equilibrium and the simulation loop is restarted. This loop continues, typically for ten or more iterations until a converged equilibrium solution is obtained. An entire simulation of this type takes between 12 to 24 hours of wall-clock time running on two nodes on the NERSC Perlmutter supercomputer. An example of an IPS workflow similar to the one described here is shown in Figure 3.
While not included in the simulations presented here, a number of additional codes are under development by Realta Fusion and collaborators which are in varying stages of IPS integration. In particular, significant progress has been made with integration of Hybrid-VPIC (Le et al., 2023) simulations into the IPS workflow such that ion-scale kinetic stability and MHD stability of the plasma equilibria computed by the IPS workflow can be evaluated (the CQL3D-m distribution functions and the equilibrium magnetic fields computed by Pleiades are used to initialize a Hybrid-VPIC simulation in these cases). A publication on this work is currently in preparation. In principle, if the spatial and velocity diffusion computed by Hybrid-VPIC can be mapped back into CQL3D-m using the IPS, it could provide a full classical and kinetic transport model for a simple mirror (as well as an assessment of MHD stability). This method of integrated transport modelling is similar to the integrated transport models used in tokamaks such as TGYRO/NEO/CGYRO (Belli & Candy, 2008, 2011; Candy et al., 2009, 2016). Edge neutrals modelling is also not included in the present model. Ingress of neutrals into the mirror plasma and their charge exchange with mirror trapped fast ions can be an important loss mechanism in small magnetic mirror experiments, and it will be necessary to include this effect to achieve quantitative simulation agreement in WHAM. Also underway is tighter integration of the IPS driven RealTwin simulations with Realta Fusion’s engineering design models. This integration will include features such as checks on component interference during IPS simulation setup and the ability to import IPS driven simulation results to form fusion neutron sources in neutronics simulations.
3 Identification of an optimized end plug with tandem mirror POPCONs
Designing an optimized end plug requires the determination of what end plug conditions produce a tandem mirror central cell capable of high gain. To do this, a modified version of the Plasma OPeration CONtour (POPCON) technique (Cordey et al., 1981; Houlberg et al., 1982) was employed. POPCONs allow the steady-state operating space of a fusion device to be mapped by enforcing a zero-dimensional power and particle balance. POPCONs have been used successfully in a number of tokamak designs, oftentimes reproducing more sophisticated transport analyses surprisingly well (Sorbom et al., 2015; Creely et al., 2020; Frank et al., 2022b). This section develops a novel version of the POPCON technique for tandem mirror central cell plasmas that, given a realistic set of end plug parameters, can be used to estimate the tandem mirror central cell performance. This POPCON technique is remarkable in that it implies that central cell performance in a classical tandem is dependent on only the end plug density and the parameters of the plug mirror coils. The POPCONs will be used here to identify a set of target end plug parameters that will be demonstrated as achievable in the RealTwin simulations in the next section.
3.1 POPCON for tandem mirrors
To generate tandem mirror POPCONs, the steady-state 0-D tandem mirror transport equations based on those found in Mirin et al. (1983); Hua & Fowler (2004) are taken to the steady state limit and all derivatives are taken to zero. However, consistent with the typical POPCON formulation, radial variation in plasma parameters (including radial variation in axial confinement) has been allowed. It is assumed that the tandem mirror is operating a classical tandem without thermal barriers (Fowler & Logan, 1977). Under this assumption electrons in the central cell and end plug are in thermal contact and $T_{e}\equiv T_{ec}=T_{ep}$. This conservative assumption dramatically simplifies the system of equations which must be solved to make a POPCON. Under these assumptions, the primary energy and particle balance equations in the POPCON system are:
$\underline{\text{Density}}$
[…]
Ion $\underline{\text{Ion Energy}}$
[…]
Electron $\underline{\text{Electron Energy}}$
[…]
where subscripts $c$ and $p$ denote central cell and end plug quantities respectively, subscripts $i$ and $e$ denote electrons and ions respectively, $n_{c}\equiv n_{ec}=n_{ic}$ is the density assuming quasineutrality, $T_{s}$ is the temperature of species $s$, $\ell_{c}$ is the central cell length, $a_{c}$ is the central cell radius, $\Gamma_{\text{gas}}$ is the gas fueling rate, $\tau_{c}\approx\tau_{p}\approx\tau_{E}$ is the confinement time, $\mathcal{P}_{ie}$ is the collisional power density transfer from ions to electrons, $P_{\mathrm{RF},s}$ is the applied RF heating power to the central cell, $\mathcal{P}_{\mathrm{fus},s}$ is the fusion power density from alpha slowing down applied to species $s$, and $\mathcal{P}_{\mathrm{rad}}$ is the power density radiated by bremsstrahlung and synchrotron radiation.
In order to create a set of POPCONs, (13) is solved for heating $P_{\mathrm{RF}}$ (with some fraction directed at ions and electrons respectively) and fueling $\Gamma_{\mathrm{gas}}$ at a given operating point with fixed $T_{ic}$ and $n_{c}$. To solve this system, all variables must be written as a function of $T_{ic}$ and $n_{c}$ and fixed end plug and central cell parameters: $B_{m}$, $a_{m}$, $n_{p}$, $\ell_{c}$, $B_{0c}$ ($T_{e}$ may be solved for as a dependent variable). This is done by applying the following definitions; the confinement time due to classical radial and axial transport (the dominant transport modes expected in the tandem mirror) $\tau_{c}$ is defined:
$\tau_{c}=\left(\frac{1}{\tau_{\text{Past}}+\tau_{f}}+\frac{1}{\tau_{\rho}}\right)^{-1},$ (14)
where the Pastukhov confinement of ions in the central cell by the end plug potential is defined (Pastukhov, 1974; Cohen et al., 1978):
$\tau_{\text{Past}}=\frac{\sqrt{\pi}}{2}\tau_{ii}\frac{\phi_{i}}{T_{ic}}\exp\left(\frac{\phi_{i}}{T_{ic}}\right)\frac{G(R_{mc})}{1+T_{i}/2\phi_{i}-(T_{ic}/2\phi_{i})^{2}},$ (15)
with $\tau_{ii}$ the central cell ion-ion pitch-angle scattering time and $\phi_{i}$ the ion confining potential between the central cell and the end plug resultant from their density difference:
$\phi_{i}=T_{e}\ln\left(\frac{n_{p}}{n_{c}}\right),$ (16)
and function $G(R_{mc})$ of the central cell mirror ratio $R_{mc}=B_{m}/(B_{0c}\sqrt{1-\beta_{c}})$ defined:
$G(x)=\sqrt{1+x^{-1}}\ln\left(\frac{\sqrt{1+x^{-1}}+1}{\sqrt{1+x^{-1}}-1}\right).$ (17)
The confinement due to collisional trapping of ions, which can become important for large $\ell_{c}$, is defined (Rognlien & Cutler, 1980):
$\tau_{f}=\sqrt{\pi}R_{mc}\frac{\ell_{c}}{v_{\mathrm{th}ic}}\exp\left(\frac{\phi_{i}}{T_{ic}}\right),$ (18)
where $v_{\mathrm{th}ic}=\sqrt{T_{ic}/2m_{ic}}$ is the ion thermal velocity in the central cell. The confinement time associated with classical radial transport is defined:
$\tau_{\rho}=\left(\frac{a_{c}}{\rho_{ic}}\right)^{2}\tau_{ii},$ (19)
where $\rho_{ic}=v_{\mathrm{th}ic}/\Omega_{ic}$ is the central cell ion Larmor radius. Returning to (13), the electron confining potential in the plug can be defined for classical tandem operation with electron-drag dominated end plugs by solving transcendental equation (Fowler & Logan, 1977; Hua & Fowler, 2004):
$\frac{\phi_{e}}{T_{e}}\exp\left(\phi_{e}/T_{e}\right)\approx\sqrt{\frac{m_{i}}{m_{e}}}\left(\frac{T_{ic}}{T_{e}}\right)^{3/2}\left(\frac{\phi_{i}}{T_{ic}}\right)\exp(\phi_{i}/T_{ic}).$ (20)
The power density associated with ion to electron energy transfer is:
$\mathcal{P}_{ie}=-\mathcal{P}_{ei}=3\frac{m_{e}}{m_{i}}\frac{n_{c}}{\tau_{ei}}\left(T_{ic}-T_{e}\right),$ (21)
where $\tau_{ei}$ is the ion-electron collision time. The fusion power density is calculated for a D-T plasma with alphas depositing their energy on species $s$ is calculated with:
$\mathcal{P}_{\text{fus},s}=f_{\alpha s}(x)\frac{E_{\alpha}}{4}n_{c}^{2}\langle\sigma v\rangle,$ (22)
using the reactivity coefficients $\langle\sigma v\rangle$ from Bosch & Hale (1992) and equation for the fraction of power deposited on each species $s$ by alpha particles from the analytic solution to the slowing down equation (Ott, 1997):
$f_{\alpha e}(x)=x^{-2}\left\{x^{2}+\frac{1}{3}\ln\left[\frac{(x+1)^{2}}{x^{2}-x+1}\right]-\frac{2}{\sqrt{3}}\tan^{-1}\left[\frac{1}{\sqrt{3}}(2x-1)\right]-\frac{\pi}{3\sqrt{3}}\right\}$ (23a)
$f_{\alpha i}(x)=1-f_{\alpha e}(x),$ (23b)
where $x=\sqrt{E_{\alpha}/E_{\text{crit}}}$ and the "critical" energy where the ion and electron energy deposition by the $\alpha$ particles is balanced $E_{\text{crit}}=[3\sqrt{\pi}/(4m_{i}\sqrt{m_{e}})]^{2/3}m_{\alpha}T_{e}=59.2T_{e}/\mu^{2/3}$ where $\mu=m_{i}/m_{p}$. Finally, the radiated power density is the sum of the power densities from bremsstrahlung and synchrotron radiation. The radiated power densities, using the expression for synchrotron radiation assuming an infinite straight cylinder with finite wall reflectivity, are (Trubnikov, 1979):
$\mathcal{P}_{\mathrm{rad}}=\mathcal{P}_{\mathrm{synch}}+\mathcal{P}_{\mathrm{brem}}$ (24a)
$\mathcal{P}_{\mathrm{brem}}=5.34\times 10^{-32}n_{e}^{2}\sqrt{T_{e}}~{}[\mathrm{W}\,\mathrm{m}^{-3}]$ (24b)
$\mathcal{P}_{\mathrm{synch}}=4.14\times 10^{-5}\sqrt{(1-R_{w})\frac{n_{e}}{a_{c}}}\left(1+2.5\frac{T_{e}}{m_{e}c^{2}}\right)(T_{e}B_{0c}\sqrt{1-\beta_{c}})^{2.5}~{}[\mathrm{W}\,\mathrm{m}^{-3}],$ (24c)
where $R_{w}$ is the wall reflectivity (taken to be 80%, the same value used TGYRO tokamak transport simulations (Candy et al., 2009)) and units of $T_{e}$ and $B_{0c}$ are taken to be in $\mathrm{keV}$ and $\mathrm{T}$ respectively.
Using this POPCON power-balance, it was possible to reproduce the tandem mirror operating points identified in Fowler et al. (2017) when the same assumptions were made i.e. no classical transport (19), fusion heating (22), or radiation (24), and all central cell RF heating was taken to be electron directed. The $T_{ic}=150~{}\mathrm{keV}$, $\beta_{c}\sim 1$ operating points described in Fowler et al. (2017) were found to within a small factor of error (<1.25) in $T_{i}$ that was determined to be associated with a discrepancy in the fusion cross sections used in their model and the Bosch-Hale cross sections used here, they considered $T_{ic}\equiv T_{ec}$ when in reality $T_{ic}\sim T_{ec}$ at these parameters, as well as the sometimes poor convergence of the model here at very high values of $\beta_{c}$. When the additional terms used here were included it was found that the operating points in Fowler et al. (2017) could not be attained. Radiation and classical transport significantly reduced performance (both terms become large at the very high $\beta\sim 1$ operating points considered there, though, it should be noted that the expression for classical transport used here is not valid when $\beta\sim 1$).
3.2 Identifying operating points with model end plugs
$a_{m}$ [m] $B_{m}$ [T] $B_{0c}$ [T] $\ell_{c}$ [m] $\langle n_{p}\rangle$ [m3]
0.10 25.0 5.0 50.0 $2.0\times 10^{20}$
0.15 25.0 3.0 50.0 $1.5\times 10^{20}$
0.20 25.0 2.6 50.0 $1.25\times 10^{20}$
0.25 25.0 2.4 50.0 $1.0\times 10^{20}$
Table 1: Tandem mirror parameters used in the POPCONs which appear in Figure04
Figure 4: POPCON plots for tandem mirrors at four different plasma radii through the mirror throat, (a) 0.1 m, (b) 0.15 m, (c) 0.2 m, and (d) 0.25 m. The other parameters used are shown in Table 1. Operating points are shown as a function of central cell density $n_{c}$ divided by end plug density $n_{p}$ versus central cell ion temperature $T_{i}$. Blue and red filled contours are of heating power to the central cell required at a given ($n_{c}$, $T_{i}$) operating point (blue regions indicate ignition). Regions which are off the scale for positive or negative and are inaccessible are denoted with red and blue hatching respectively. Black contour lines are of fusion power in the central cell. Green contour lines are of central cell vacuum $\beta$.
With the components of the POPCON system of equations (13) defined, it is possible to solve for the required heating and fueling at each location in ($n_{c}$, $T_{ic}$) space for a given set of end plug radial density profiles and engineering parameters. POPCONs with simplified end plugs were solved to establish target parameters for the end plug to be simulated with RealTwin. In these simplified end plugs, the density profile was taken to be flat, the ion directed fraction of $P_{\mathrm{RF}}$ was taken to be $80\%$, and the end plug and central cell parameters shown in Table 1 were used. End plug parameters were taken to be fixed besides the plasma radius through the mirror throat $a_{m}$ and the average density within the plug $\langle n\rangle$ which were varied to determine a target operating range. The central field was increased until either the number of alpha gyroradii across the plasma radius $N_{\alpha}>5$ defined using:
$N_{\alpha}=\frac{a_{0c}}{\rho_{\alpha}}\approx 5.24a_{m}\sqrt{B_{m}B_{0c}}$ (25)
was obtained or an ignited operating point with $P_{\mathrm{fus}}>150~{}\mathrm{MW}$ and $\beta<0.6$ was obtained. Average end plug density $\langle n_{p}\rangle$ was varied between $1.0\times 10^{20}~{}\mathrm{m}^{-3}$ and $2.0\times 10^{20}~{}\mathrm{m}^{-3}$.
The results of the POPCON analysis described here are shown in Figure 4 for the parameters defined in Table 1. POPCON analysis indicates that, for what will be shown later on are fairly reasonable end plug parameters, it is possible to create a classical tandem mirror with an ignited central cell. Large ignition regions are present in all cases when the plasma radius at the mirror throat is larger than 0.15 m. These results indicate that a tandem mirror pilot plant designed with these principles should easily be able to satisfy the NASEM fusion pilot plant requirements at lengths under 100 m. Using simplified equation for the net electric power:
$P_{\mathrm{net,elec}}=\left(P_{\mathrm{fus},\alpha}+P_{\mathrm{NBI}}+C_{\mathrm{mult}}P_{\mathrm{fus},n}\right)\eta_{\mathrm{ele}}-\frac{P_{\mathrm{NBI}}}{\eta_{\mathrm{NBI}}}$ (26)
Assuming neutral beam efficiencies $\eta_{\mathrm{NBI}}=60\%$, an electrical conversion efficiency $\eta_{\mathrm{ele}}=50\%$ assuming the use of a Brayton cycle (Schleicher et al., 2001), a blanket neutron energy multiplication factor $C_{\mathrm{mult}}=1.1$, and $P_{\mathrm{NBI}}=30~{}\mathrm{MW}$ into the end plugs; it was found a tandem mirror would need to produce $157.4~{}\mathrm{MW}$ of power to satisfy the NASEM $50~{}\mathrm{MWe}$ requirement. Such an operating point is achievable with $B_{m}=25~{}\mathrm{T}$, $a_{m}=0.15~{}\mathrm{m}$, $\langle n\rangle_{p}=1.5\times 10^{20}~{}\mathrm{m}^{-3}$, $\ell_{c}=50~{}\mathrm{m}$, $\beta_{c}\cong 0.6$, $n_{c}/n_{p}\cong 0.55$, $T_{i}\cong 45~{}\mathrm{keV}$, and $T_{e}\cong 125~{}\mathrm{keV}$ based on this POPCON analysis. The fusion power density of this case is $\sim 3.5~{}\mathrm{MWm}^{-3}$ which is competitive with high-field tokamaks (Frank et al., 2022b). Performance of such a system could be further enhanced with technologies such as direct electrical conversion (Moir & Barr, 1973) or more efficient NBI utilizing advanced neutralization techniques and energy recovery (Fumelli et al., 1989; Grisham, 2009; Kovari & Crowley, 2010). It is noteworthy that this level of performance is achievable despite using generally more conservative parameters than the DT classical tandem proposed in Fowler et al. (2017) which had: $B_{m}=24~{}\mathrm{T}$, $a_{m}=0.21~{}\mathrm{m}$, $\langle n_{p}\rangle=2.6\times 10^{20}~{}\mathrm{m}^{-3}$, $\ell_{c}=55~{}\mathrm{m}$, $\beta_{c}\sim 1.0$, $n_{c}/n_{p}\sim 0.35$, $T_{i}=150~{}\mathrm{keV}$, and $T_{e}=150~{}\mathrm{keV}$.
Figure 5: Contours of $T_{e}$ for operating points at given $\langle n\rangle$ and $E_{\mathrm{NBI}}$ satisfying the constraint equations (34). This plot uses the $a_{m}=0.15$ m POPCONs case found in Figure 4b.
This POPCONs analysis exposed a number of surprising features of fusion relevant tandem mirror operation that were not previously well known. These discoveries are in part because, outside of the analysis in Fowler et al. (2017), axisymmetric tandem mirrors with high performance HTS magnets and high $R_{m}$ had never been considered. Older studies of tandem mirrors generally assumed end plug mirror ratios of order unity to account for minimum-B coil geometries and lower peak $B_{m}$ (Fowler & Logan, 1977). This analysis builds upon that in Fowler et al. (2017), and very importantly, it includes the effect of fusion alpha particle heating on the electron temperature. The key result here is that, when high-field mirror coils are available, classical tandem central cells can and should be operated at much higher density fractions $n_{c}/n_{p}$ than previously assumed. Much of the previous classical tandem mirror literature used $n_{c}/n_{p}=0.05-0.35$ (Dimov et al., 1976; Fowler & Logan, 1977; Hua & Fowler, 2004; Fowler et al., 2017). The analysis here suggests this value should be higher, $n_{c}/n_{p}=0.35-0.6$. Strong alpha heating of the electrons obtained by, increasing the central cell density, boosting confinement with large $R_{m}$ from axisymmetrization, and directly heating the central cell ions with RF to enhance the fusion rate, allows the central cell to reach electron temperatures $\sim 2-3~{}T_{i}$ without supplemental electron heating as shown in Figure 5. This operational mode takes advantage of tandem mirror confinement (15) improving more rapidly with increasing temperature than it degrades with increasing the density ratio.
In conclusion, this scoping shows that for high-field axisymmetric mirrors there classical tandems are a valid alternative to the traditional thermal-barriered tandem operation (Baldwin & Logan, 1979) which can still reach appreciable Q in agreement with similar analyses by Pratt & Horton (2006) and Fowler et al. (2017). However, it has been found here that higher density operation than these previous analyses is possible thanks to alpha heating, somewhat relaxing the end plug performance criteria, but there are still some important caveats. The tandem mirror POPCON approach used here, much like the tokamak POPCON approach, is only able to identify operating points. It does not guarantee that an operating point it identifies is stable or achievable in actual operation. To determine this, a time dependent simulation which takes the mirror from an initial breakdown plasma to its final operating point must be performed. It is possible that, to actually reach the operational conditions proposed here, a transient heating source, such as a high-power electron cyclotron heating "sparkplug", might be needed to boost $T_{e}$ initially. This would drive up the end plug potentials and improve confinement enough for the central cell to generate sufficient alpha heating power to reach ignition. Furthermore, the effects of spatial gradients or turbulent transport were not included in this analysis. Generally, the spatial density gradients in tandem mirrors are predicted to be small (Mirin et al., 1983; Hua & Fowler, 2004), and the effects of electron temperature gradient turbulence, the main form of turbulent energy transport predicted to be present in tandem mirrors, are anticipated to be more tolerable than the temperature gradient driven turbulent transport in tokamaks (Hua & Fowler, 2004; Pratt & Horton, 2006). These topics will be treated in future work where gyrokinetic turbulence simulations at mirror central cell conditions will be performed to evaluate the growth rates of gradient driven instabilities. A time dependent solver based on those found in Mirin et al. (1983) and Hua & Fowler (2004) in order to evaluate the accessibility of these operating points under more realistic assumptions will also be implemented. In addition to these confinement issues, macrostability of the tandem system, specifically its stability against trapped particle modes and MHD, have not been addressed here and will be the topic of a future publication.
One potentially interesting feature of the relatively high $n_{c}/n_{p}$ ratio used in this study is the question of whether outflow from the central cell may have a stabilizing affect on kinetic modes in the end plugs. The flux from the central cell through the end plug midplane can be determined by particle balance of the outflow along the flux tube:
$\Gamma_{c,\mathrm{out}}=\frac{n_{c}\ell_{c}}{2\tau_{c}}\left(\frac{R_{mc}}{R_{mp}}\right)$ (27)
If this outflow flux becomes large enough, it has been shown to be stabilizing to DCLC (Drake et al., 1981), but large amounts of central cell outflow could also reduce an end plug’s classical performance by increasing the ion-ion collisional scattering rate. Taking the estimate for the outflow required to stabilize the plug from Correll et al. (1980):
$\Gamma_{p,\mathrm{stab}}>1.2\times 10^{5}\frac{n_{p}\phi_{e}^{3/2}}{(R_{mp}\sin^{2}\theta_{\mathrm{NBI}}-1)\sqrt{\mu}E_{ip}}\left(\frac{a_{m}\sqrt{R_{mp}}}{\rho_{i0}}\right)^{-4/3}~{}[\mathrm{m}^{-2}\mathrm{s}^{-1}],$ (28)
where $\theta_{\mathrm{NBI}}$ is the angle of the NBI in the plug, $\mu$ is the average ion mass in the plug in AMU, $\phi_{e}$ is the plug potential in kV, $E_{\mathrm{NBI}}$ is the plug NBI energy in keV (all other quantities are assumed to be in MKS), and $\rho_{i0}$ is ion Larmor radius at the midplane of the plug. Using these two equations, the minimum $n_{c}/n_{p}$ density ratio at which DCLC stabilization in the plug can be obtained:
$\frac{n_{c}}{n_{p}}>2.4\times 10^{5}\frac{\tau_{c}\phi_{e}^{3/2}}{\ell_{c}(R_{mp}\sin^{2}\theta_{\mathrm{NBI}}-1)\sqrt{\mu}E_{ip}}\left(\frac{R_{mp}}{R_{mc}}\right)\left(\frac{a_{m}\sqrt{R_{mp}}}{\rho_{i0}}\right)^{-4/3}.$ (29)
For tandem mirror parameters similar to the ignited operating points in Figure 4b: $\tau_{c}=5~{}\mathrm{s}$, $\phi_{e}=400~{}\mathrm{keV}$, $\ell_{c}=50~{}\mathrm{m}$, $R_{mp}=8$, $R_{mc}=13.3$ (accounting for finite $\beta\sim 0.6$ in both mirror ratios), $\theta_{\mathrm{NBI}}=45^{\circ}$, $E_{i}=500~{}\mathrm{keV}$, $\mu=2.5~{}\mathrm{AMU}$, and $a_{m}\sqrt{R_{mp}}/\rho_{i0}=25$, it is calculated that $n_{c}/n_{p}>1000$ indicating DCLC likely cannot be stabilized with central cell outflow in the tandem mirrors under consideration here. As $\Gamma_{c,\mathrm{out}}$ is insufficient for DCLC stabilization, DCLC stability in Section 4 will be ensured by the suppression of the diamagnetic drift wave and the use of sloshing ions.
Finally, there is the matter of alpha particle removal from the central cell in steady-state operation. Alpha particles are well confined in the central cell by the end plug’s ambipolar potential. Without a method to deconfine them, they will accumulate. This will need to be addressed with a technique such as drift pumping, which removes thermalized alphas by applying a radiofrequency perturbation to the background magnetic field that causes them to drift out of the plasma after enough central cell transits (Dimov et al., 1976; Logan, 1983).
4 Simulation of an optimized end plug using RealTwin
The POPCON technique outlined in the previous section has defined what parameters the end plug must achieve in order to obtain a good central cell: the end plug mirror coils must have the highest field and bore possible, and an end plug must achieve the highest plasma density possible for the fewest watts of input power. In this section an end plug will be optimized, first some assumptions will be used to simplify the problem and physics and engineering constraints will be imposed on the end plug. Next, those constraints will be used to perform an initial optimization of the plug and identify an operating region. Finally, simulations with the RealTwin will be performed to provide a high-fidelity optimization based on the prior constraint based optimization.
4.1 Assumptions and constraints
When performing simulations with the RealTwin model, several assumptions about end plug operation were made due to the limitations of present mirror modelling tools (a discussion of ways in which these capability gaps may be closed in the future appears in Section 5). In addition, a number of constraints related to engineering feasibility and plasma stability were used to do significant optimization work prior to even running the RealTwin.
Beginning with the modelling assumptions and fixed parameters; two large physics assumptions were made in this study to overcome the limitations of present modelling tools. First, it was assumed that the end plug, for the purposes of equilibrium and transport modelling, could be treated as a standalone simple mirror. Second, the question of MHD stabilization of the end plug was ignored. The justification of the first assumption lies in the fact that the ion populations in the plug and the central cell are to lowest order "detached". Ions confined in the central cell do not enter the plugs and ions confined in the plugs do not enter the central cell. Similarly ions that are not confined in the plug are also not confined in the central cell and vice-versa. This property of tandem mirrors ends up being useful for reasons that will be discussed momentarily. In fusion-relevant conditions, confinement is good in both the plugs and the central cell and the unconfined population of ions flowing through either region is small, acting as a (relatively) low density stream which passes outwards until it reaches the expander region (Cohen et al., 1980). However, in a classical tandem mirror the electron population is able to transit between the plug and central cell, and alpha particle heating from the central cell is sufficient to enable us to reach high electron temperatures. In fact, the predicted electron temperatures in a tandem are higher than they otherwise would be in a standalone end plug. This makes the estimates of plug performance in this model conservative with respect to electron drag, as the increasing electron temperature will decrease electron drag driven ion losses and improve ion confinement (however there can be increased ion losses due to larger ambipolar potentials which may incentivize using higher neutral beam energies). A more detailed analysis of how varying $T_{e}$ affects ion confinement can be found in Appendix A. For the purposes of the simulations here, it was assumed that the plasma is MHD stable and whichever actuator used to ensure MHD stability was assumed to not substantially perturb equilibrium transport. This topic will be tackled in more detail in future work.
As mentioned earlier, the detached nature of the end plug and central cell ion distributions can be used to improve performance. In a tandem, the end plug ion composition need not be the same as the central cell as the end plug and central cell ion populations are not exchanged (Fowler & Logan, 1977; Cohen et al., 1980). This presents a unique opportunity to enhance end plug physics performance and simplify end plug engineering. To this end, end plugs fueled with only tritium are used in conjunction with a deuterium and tritium fueled central cell. This combination maximizes end plug performance and reduces the nuclear engineering challenges presented by the end plug. The physical argument for this operational mode is as follows: ion confinement in the end plug is dominated by two processes: ion-ion scattering and electron drag. The particle confinement times associated with these two processes are (Logan et al., 1980):
$\tau_{s}$ $=2.76\times 10^{18}\mu^{1/2}E_{p}^{3/2}\frac{\log_{10}(R_{m}\sin^{2}\theta_{\text{NBI}})}{Z^{4}n_{i}\ln\Lambda_{ii}}$ (30a)
$\tau_{d}$ $=1\times 10^{14}\mu T_{e,p}^{3/2}\frac{\ln(E_{\mathrm{NBI}}/\langle E_{L}\rangle)}{Z^{2}n_{e}\ln\Lambda_{ei}}$ (30b)
$\tau_{p}$ $\approx\left(1/\tau_{s}+1/\tau_{d}\right)^{-1},$ (30c)
where the average ion loss energy $E_{L}$ can be defined:
$E_{L}\approx\frac{1}{1+(\tau_{s}/\tau_{d})}\left[E_{\mathrm{NBI}}+E_{h}\left(\frac{\tau_{s}}{\tau_{d}}\right)\right],$ (31)
with ambipolar hole loss energy:
$E_{h}=\frac{\phi_{e}}{R_{m}\sin^{2}\theta_{\text{NBI}}-1}.$ (32)
At typical tandem operational parameters, confinement is assumed to be electron drag dominated, $\tau_{d}\ll\tau_{s}$ (Logan et al., 1980; Mirin et al., 1983; Hua & Fowler, 2004; Fowler et al., 2017). As the end plug ion population composition need not be the same as the central cell, performance can be enhanced by choosing ions with large $\mu/Z^{2}$. In addition, choosing an ion mix to suppress fusion neutron rates in the end plug can substantially simplify engineering by reducing radiation shielding requirements for components such as magnets and NBI. Suppressing fusion rates in the end plugs is acceptable as the total fusion power produced by the end plug is a small component of the overall fusion power in a tandem, since the end plug volume is much less than central cell volume, $V_{p}\ll V_{c}$. If taken to its logical conclusion, this argument suggests that a tritium (T) only end plug is optimal with respect to classical transport and neutron damage. Tritium has the highest $\mu/Z^{2}$ ratio of any standard fusion fuel ion, and the T-T fusion cross-section is nearly an order of magnitude lower than the D-D fusion cross section, and T-T fusion produces neutrons with energies of $\sim 3~{}\mathrm{MeV}$ (Brown et al., 2018). Thus, a T-only end plug provides a factor of $\sim 1.2$ increase in confinement performance and a many-order of magnitude reduction in the neutron rate as well as a significant reduction in neutron energy relative to a D-T end plug (T fueled plugs will have inferior microstability to D or H plugs as the T Larmor radius is larger, making it easier to excite diamagnetic drift waves. However, it will be shown that it should be possible to avoid diamagnetic drift waves with large $a/\rho_{i}$ even with T plugs).
Another assumption is the use of only negative-ion neutral beams for end plug heating in steady-state (a discussion about the limitations associated with high-harmonic RF heating coupled with positive-ion NBI, similar to what is planned in WHAM, and the reason it is not used in this work is discussed in Appendix B). The use of ECH has also been limited to achieving breakdown (assuming a scheme like that found in Yakovlev et al. (2016)). Presently, the use of any significant quantity of ECH in CQL3D-m causes the potential profile to invert. This behavior is actually promising as it indicates that it may be possible to generate thermal barriers (Baldwin & Logan, 1979), but at the moment it causes significant problems with the bounce-averaging and ambipolar potential calculations in CQL3D-m. Thus, further investigation of ECH is left for future work. Neutral transport into the tandem mirror plasma has also been neglected. This effect is anticipated to be of some importance in small relatively low density mirrors such as WHAM, where neutrals recycled from the edge can enter the plasma and charge exchange with well-confined beam ions causing them to be lost and replacing them with a promptly lost cold ion. However, fusion-grade plasmas like those under consideration here are opaque to edge neutrals, and the effect of these neutrals on transport can be ignored. The final assumption made here is that transport in the end plug will remain classical during operation. To this end, it has been insured that the DCLC stability condition is well satisfied in the plasmas optimized here.
Figure 6: An example of the magnetic configuration used in the RealTwin end plug simulations. The high-field (HF) mirror coils are shown in red and the low-field (LF) central coil is shown in black. Contours of poloidal flux are shown in blue and the limiting surface is shown with a black dashed line.
In addition to physics assumptions, some engineering constraints were imposed on the mirror design here. First, the field of the two large bore high-field mirror coils placed at either end of the end plug (denoted HF) was fixed at $25~{}\mathrm{T}$. Feasibility studies of HTS coils at these fields have found that such designs should be viable at bores sizes appropriate for fusion applications, $\sim 1$ m warm bore diameter. A single low-field coil (denoted LF) was placed at the center of the end plug to give control of the central field. The magnetic configuration for an example case is shown in Figure 6. The minimum length $\ell$ (where subscript $p$ on end plug quantities in this section has been dropped for simplicity), defined as the mirror throat to mirror throat distance. This limit was chosen not by the finite Larmor radius (FLR) stabilization condition for $m\geq 2$ modes, but rather by engineering constraints. The substantial size of the neutral curvature central cell in addition to the good curvature expanders is expected to provide significant stabilization against high $m$ modes (Roberts & Taylor, 1962; Rosenbluth et al., 1962; Rosenbluth, 1965; Pearlstein, 1966; Anikeev et al., 1992; Ryutov et al., 2011; Fowler et al., 2017). Active stabilization schemes are anticipated to provide for remaining MHD stabilization needs. Here, $\ell\geq 4.5~{}\mathrm{m}$ has been chosen as with values of $\ell<4.5~{}\mathrm{m}$ require large reverse polarity currents in the LF coil to cancel out the HF field. Additionally, very short end plugs will almost certainly impede NBI port access and require the use of inefficient NBI launch angles that limit beam path length, causing reduced NBI absorption. The optimal NBI configuration for obtaining sloshing ion distributions, maximizing beam path length, and maintaining good confinement (noting there is an effective mirror ratio reduction associated with shallower NBI launch angles) occurs at, or very close to, a neutral beam with $\theta_{\mathrm{NBI}}=45^{\circ}$ targeted at the midplane. To maintain an effective $45^{\circ}$ ion velocity-space angle and obtain a sloshing ions distribution for different beam launch angles, the beam’s target location must be moved to different values of $B/B_{0}$. The reduction factor in beam path length incurred by this move is in a simple mirror is $\sqrt{(1+\cot^{2}\theta_{\mathrm{NBI}})/4}/\sin\theta_{\mathrm{NBI}}$. To simplify this study, only the near optimal case of $\theta_{\mathrm{NBI}}=45^{\circ}$ with the beam directed at the midplane was examined in detail (the NBI here used a $0.4~{}\mathrm{m}$ by $0.8~{}\mathrm{m}$ source with no beam divergence as a placeholder until more realistic NBI specifications are known). In later work, with better defined, beam, magnet, and radiation shielding parameters, this angle may be slightly adjusted to prevent beam-line interference with components.
Figure 7: Contours of $a_{m}$ (red lines) and $P_{\mathrm{NBI}}$ (filled contours) for operating points at given $\langle n\rangle$ and $E_{\mathrm{NBI}}$ satisfying the constraint equations (34). The grey region represents operation regimes requiring greater than 20 MW of NBI power to access.
After fixing some engineering parameters, a basic physics constraint based optimization was used to identify a parameter space which seemed likely to produce attractive results. To do this, the primary parameters of an negative-ion NBI driven plug: $a_{m}$, $\ell$, $B_{m}$, $B_{0}$, $E_{\mathrm{NBI}}$, $\langle n\rangle$, $B_{0}$, and $P_{\mathrm{NBI}}$ were defined. From these parameters, most other major parameters in the plasma can be defined. Fixed parameters were $B_{m}=25~{}\mathrm{T}$ and $\ell=4.5~{}\mathrm{m}$. From the POPCON analysis in Section 3, the acceptable values of $\langle n\rangle$ were known to range from $1\times 10^{20}~{}\mathrm{m}^{-3}$ to $2\times 10^{20}~{}\mathrm{m}^{-3}$. By fixing $\langle n\rangle$ and $E_{\mathrm{NBI}}\sim E_{i}$ to some desired values, it is possible to presume that the electron energy is $\sim 0.1E_{\mathrm{NBI}}$ which Egedal et al. (2022); Endrizzi et al. (2023) have shown is roughly accurate for an NBI heated simple mirror plasma plasma. The beam path was estimated to be $\sim 2\sqrt{2}a_{m}$ and the plug volume was taken to be $V\sim 2/3\pi R_{m}a_{m}^{2}\ell$ where the factor of 2/3 is added based on simulation results to account for the field curvature reducing the volume from a simple cylinder. With these assumptions in place, it was possible to constrain the rest of the mirror’s parameters by simultaneously solving system:
DCLC $\underline{\text{DCLC Stability}}$
$\frac{a_{0}}{\rho_{i}}\geq 25$ (33a)
NBI $\underline{\text{NBI Absorption}}$
[…]
for $a_{m}$ and $B_{0}$, where (33a) is a slightly modified version of the the drift-cyclotron wave stability criterion which appears in Baldwin (1977) (if the drift-cyclotron instability is stable the DCLC will not grow), and (33b) uses the analytic NBI stopping cross sections $\sigma_{s}$ from Janev et al. (1989). Using the assumptions outlined above, (33) may be rewritten (taking (33a) to be equal to 25):
DCLC $\underline{\text{DCLC Stability}}$
$B_{0}^{2}-0.16\left(\frac{\mu E_{i,100\mathrm{keV}}}{B_{m}a_{m}^{2}}\right)-3.0E_{i,100\mathrm{keV}}\langle n_{20}\rangle=0$ (34a)
NBI $\underline{\text{NBI Absorption}}$
$6.7\times 10^{-21}=\langle n_{20}\rangle\frac{\sqrt{1+\cot^{2}(\theta_{\mathrm{NBI}}/4)}}{\sin\theta_{\mathrm{NBI}}}a_{m}\sqrt{\frac{B_{m}}{B_{0}(1-\beta)^{1/2}}}\sigma_{s},$ (34b)
where ion energies are in 100’s of keV and densities are in $10^{20}~{}\mathrm{m}^{-3}$ (other units are standard MKS). At high values of $\langle n\rangle$ and $E_{\mathrm{NBI}}$, the constraint on $B_{0}$ set by $\beta$ rather than DCLC stability. In this case $B_{0}$ was constrained with condition $\beta\leq 0.5$. Once the constraint equations were solved, from the simple mirror confinement time scaling law $n_{20}\tau_{p}=0.25E_{\mathrm{NBI},100\mathrm{keV}}^{3/2}\log_{10}(R_{m})$ from Egedal et al. (2022) (this scaling law is for deuterium and does not include all the details of the scaling laws presented in (30), but is sufficient for the constraint exercise here), one can calculate the upper bound on the absorbed $P_{\mathrm{NBI}}$ required to reach the target value of $\langle n\rangle$ using fueling rate balance:
$P_{\mathrm{NBI}}=\frac{E_{\mathrm{NBI}}\langle n\rangle V}{\tau_{p}}$ (35)
From the POPCON analysis performed in Section 3, it was identified that plasma radius $a_{m}\gtrsim 0.15~{}\mathrm{m}$ and density $\langle n\rangle\sim$1.0-1.5×$\times 10^{20}~{}\mathrm{m}^{-3}$. Plotting contours of $P_{\mathrm{NBI}}$ and $a_{m}$ versus $E_{\mathrm{NBI}}$ and $\langle n\rangle$ produces a plot like that shown in Figure 7. This plot indicates an operating region accessible with reasonable $P_{\mathrm{NBI}}$ values $<25~{}\mathrm{MW}$ exists between $E_{\mathrm{NBI}}$ of $200~{}\mathrm{keV}$ and $300~{}\mathrm{keV}$. However, the method used to constrain NBI absorption with (33b) is perhaps suboptimal. It has not accounted for that, despite not being absorbed as readily at high energies, the beam may still be more efficient at fueling, and simplified scaling laws were used here rather than an actual Fokker-Planck calculation. This will be corrected in the next part of this section with more detailed simulations using the RealTwin.
4.2 Parametric simulations of end plug performance
Parameter Range
$a_{m}$ 0.15 - 0.25 m
$\ell_{p}$ 4.5 m
$B_{m}$ 25 T
$B_{0}$ 4.0 - 6.0 T
$E_{\mathrm{NBI}}$ 240 - 480 keV
$P_{\mathrm{NBI}}$ 10 - 20 MW
Table 2: End plug parameters used in the parametric scan
With the constraints in the previous section established, a parametric scan in $B_{0}$, $P_{\mathrm{NBI}}$, $E_{\mathrm{NBI}}$ and $a_{m}$ was carried out with the RealTwin model described in Section 2 was specified. This scan used the end plug parameters shown in Table 2. For this work, points in the scan were selected over a linear range for each of the parameters. This scoping method is coarse, but substantial constraints were already placed on the problem through the use of the POPCON analysis in Section 3 and the analytic arguments used to constrain the physics and engineering earlier in this section. This made the coarse high fidelity scan sufficient to deliver an optimized plug that exceeded the requirements identified by the POPCON analysis. A scan in $\ell$ was also performed to determine if scaling power density with $\ell$ versus $P_{\mathrm{NBI}}$ was substantially different (changing the aspect ratio of the plug can slightly change NBI path length and modify NBI absorption). Varying $\ell$, however, had effectively the same impact on power density and plug performance as varying $P_{\mathrm{NBI}}$. In the rest of the discussion here will focus on the highest performing $\ell=4.5~{}\mathrm{m}$ case.
The CQL3D-m simulations performed in this section used 24 flux surfaces evenly spaced from 0.01 to 0.9 in normalized square-root poloidal flux $\sqrt{\psi_{n}}$, and the axial $z$ grid used to formulate the Pastukhov response along the field line had 64 evenly spaced points extending from the midplane to the mirror throat. The distribution functions were discretized over 500 points in momentum-per-rest-mass $u_{0}$ with a grid which used increased resolution at low energies as to accurate capture the ion distributions and 400 evenly spaced points in the pitch angle $\vartheta$. The maximum energy in the momentum-per-rest-mass grid was set to correspond to a $250~{}\mathrm{keV}$ electron energy. The CQL3D-m simulations used 800 timesteps. The time advance was initially set to $1~{}\mathrm{ms}$ for the first 200 time steps, then increased to $2~{}\mathrm{ms}$ for the next 200 time steps, and then further increased to $4~{}\mathrm{ms}$ for the remaining 400 timesteps providing an overall simulation runtime of $2.2~{}\mathrm{s}$, several $\tau_{p}$. Ramped time-stepping was used to maintain numerical stability, as despite being an implicit time-advance code, the iterations of CQL3D-m with Freya and Pleiades and the internal Pastukhov potential solve are explicit. The Freya NBI deposition calculation was run at each time step using $2.5\times 10^{6}$ particles, and the Pleiades equilibrium was updated after every 100 time steps.
Figure 8: Contour plots of $\langle n\rangle$ for different values of $a_{m}$ and $B_{0}$ at three different $E_{\mathrm{NBI}}$ for an end plug with fixed length $\ell=4.5~{}\mathrm{m}$ with three different NBI heating powers $10~{}\mathrm{MW}$ (top), $15~{}\mathrm{MW}$ (middle), and $20~{}\mathrm{MW}$ (bottom).
The results of the end plug simulations using the RealTwin model are shown in Figure 8. end plugs which failed to satisfy the DCLC constraint (33a) or in which the equilibrium failed to converge due to $\beta$ becoming too high were removed from this analysis (this is why there is an empty regions in the Figure 8 plots). A potentially surprising result of these scans is that there is little difference in peak end plug density versus $E_{\mathrm{NBI}}$, however, remembering the analysis in Section 4.1, this makes sense. Neutral beam absorption drops with higher $E_{\mathrm{NBI}}$, but the NBI fueling efficiency increases $\propto E_{\mathrm{NBI}}^{1/2}$. Another observation is that $\langle n\rangle$ improves superlinearly with increasing $P_{\mathrm{NBI}}$. This is due to a twofold non-linearity in the NBI absorption equation (33b) with NBI power. As $P_{\mathrm{NBI}}$ increases, the $\langle n\rangle$ increases and therefore the neutral beam absorption fraction increases. In addition, the plasma $\beta$ also increases as $P_{\mathrm{NBI}}$ increases. This causes the plasma to expand outwards from diagmagnetism, increasing the NBI absorption path length, further increasing NBI absorption.
Figure 9: Density $n_{p}$ and average ion energy $\langle E_{i}\rangle$ and electron energy $\langle E_{e}\rangle\equiv 1.5T_{e}$ versus the square root of the normalized poloidal flux $\sqrt{\psi_{n}}$. The profiles come from a simulation with parameters: $a_{m}=0.15~{}\mathrm{m}$, $B_{m}=25~{}\mathrm{T}$, $B_{0}=4~{}\mathrm{T}$, $\ell=4.5~{}\mathrm{m}$, $E_{\mathrm{NBI}}=240~{}\mathrm{keV}$, and $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$.
As anticipated in Section 4.1, the target end plug $\langle n\rangle$ values identified in Section 3 were achieved at the lowest $P_{\mathrm{NBI}}$ for $E_{\mathrm{NBI}}$ of $240~{}\mathrm{keV}$ and $360~{}\mathrm{keV}$. Sufficient end plug conditions for fusion relevant central cell operation were achievable with $B_{0}$ of $4-6~{}\mathrm{T}$, $P_{\mathrm{NBI}}\sim 15~{}\mathrm{MW}$, $E_{\mathrm{NBI}}\sim 240-360~{}\mathrm{keV}$, and $a_{m}$ of $0.15-0.2~{}\mathrm{m}$. Under these conditions effectively full NBI absorption (>95%) was obtained with a $240~{}\mathrm{keV}$ beam, so lowering NBI energy further would only serve to reduce performance. The plasma profiles of the most optimal plug identified in this scan are shown in Figure 9. The plug had parameters: $a_{m}=0.15~{}\mathrm{m}$, $B_{0}=4~{}\mathrm{T}$, $E_{\mathrm{NBI}}=240~{}\mathrm{keV}$, and $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$ and achieved $\langle n\rangle>1.5\times 10^{20}~{}\mathrm{m}^{-3}$, more than sufficient to drive a $50~{}\mathrm{m}$ long central cell to $Q>5$ conditions based on the analysis in Section 3. This plug in particular was chosen as it has a low $a_{m}$ and $E_{\mathrm{NBI}}$, simplifying engineering, while still delivering good plasma performance. Another promising, but more technologically aggressive, end plug condition that may have better DCLC stability due to its larger $a/\rho_{i}$ was obtained with $a_{m}=0.2~{}\mathrm{m}$, $B_{0}=5~{}\mathrm{T}$, $E_{\mathrm{NBI}}=360~{}\mathrm{keV}$, and $P_{\mathrm{NBI}}=20~{}\mathrm{MW}$ achieved $\langle n\rangle\sim 1.3\times 10^{20}~{}\mathrm{m}^{-3}$ also sufficient to drive a $50~{}\mathrm{m}$ central cell to $Q>5$ conditions, indicating some degree of design flexibility. The optimized tandem mirror operating points identified here are summarized in Table 3.
Parameter Optimum Alternate
End plug
$a_{m}$ 0.15 m 0.20 m
$\ell_{p}$ 4.5 m 4.5 m
$B_{m}$ 25 T 25 T
$B_{0}$ 4.0 T 5.0 T
$\langle n_{p0}\rangle$ $1.66\times 10^{20}$ m-3 $1.35\times 10^{20}$ m-3
$\langle\beta_{p0}\rangle$ 0.58 0.53
$E_{\mathrm{NBI}}$ 240 keV 360 keV
$P_{\mathrm{NBI}}$ 15 MW 20 MW
Central cell
$a_{c}$ 0.54 m 0.78 m
$l_{c}$ 50 m 50 m
$B_{0c}$ 3.0 T 2.6 T
$\beta_{c}$ 0.6 0.6
$n_{c}$ $7.5\times 10^{19}$ m-3 $5.8\times 10^{19}$ m-3
$T_{ic}$ 50 keV 57 keV
$T_{ec}$ 100 keV 120 keV
$P_{\mathrm{fus}}$ 175 MW 200 MW
$Q$ 5.8 5
Table 3: Optimized tandem parameters based on the results of parametric scans and POPCON analysis using the end plug conditions obtained in the scans.
While the optimized plugs identified in this scan are sufficient to deliver a viable central cell, this scan was coarse. It is foreseeable that performance, i.e. $\langle n\rangle/P_{\mathrm{NBI}}$, could be improved further by using Bayesian optimization coupled with a technique like Gaussian process machine learning. Use of such an approach would facilitate more efficient sampling of points in the multidimensional optimization space ($E_{\mathrm{NBI}}$, $P_{\mathrm{NBI}}$, $a_{m}$, $B_{m}$, $B_{0}$, $\ell$) such that a more optimized operating point could be found. It also could allow expansion of the optimization space to include parameters such as detailed aspects of the NBI injection geometry (focal length, source size, etc.), and optimization of the NBI beamline interference versus the real magnet geometry. The scalable nature of the IPS driven simulations here should facilitate such optimization procedures, and similar optimization procedures have been successful when applied to various aspects of tokamak design in the past (Nunn et al., 2023; Pavone et al., 2023; Brown et al., 2024; Rodriguez-Fernandez et al., 2024).
5 Conclusions and future work
Here, a new computational model for magnetic equilibrium, transport, heating, and fueling in simple mirrors was developed, a novel POPCON technique for tandem mirror performance was derived and optimized end plugs that deliver tandem mirror central cell performance consistent with the fusion pilot plant criteria outlined in the NASEM recommendations in the "Bringing Fusion to the US Grid" report were identified. The analysis here has identified an optimized $25~{}\mathrm{T}$ peak field tandem mirror end plug that achieves average densities in excess of $1.5\times 10^{20}~{}\mathrm{m}^{-3}$ with $15~{}\mathrm{MW}$ of $240-360~{}\mathrm{keV}$ NBI heating and enables a $50~{}\mathrm{m}$ long $3~{}\mathrm{T}$ central cell to generate over $150~{}\mathrm{MW}$ fusion power. This operating point is significantly more conservative than previous classical tandem mirror concepts proposed in the literature (Fowler & Logan, 1977; Fowler et al., 2017). Along the way, a number of important new and important insights about classical axisymmetric tandem mirror operation were brought to light. Using the novel POPCONs approach derived here, it was found found that in a classical tandem mirror the end plug can be characterized sufficiently to estimate central cell performance with only parameters: $a_{m}$ the plasma radius at the mirror throat, $B_{m}$ the field at the mirror throat, and $n_{p}$ the end plug density. The POPCONs found that alpha particle heating in the central cell can substantially raise $T_{e}$ and is important to understanding tandem mirror performance, and the central cell density ratio $n_{c}/n_{p}$ in high-field axisymmetric tandem mirrors should be significantly higher than previously suggested in the literature. End plug performance and engineering feasibility were also found to be enhanced by fueling the end plug with only tritium. Finally, first of a kind optimizations with integrated modelling for simple mirrors was performed using the RealTwin and it was shown that pilot plant end plug parameters can be reached.
Along with these new discoveries, this work has exposed a number of important computational and theoretical capability gaps which need to be filled to further the modelling of tandem mirrors. Most importantly, the models here used a piece-wise approach to estimate central cell performance based on end plug performance. This required a number of assumptions. Ideally, a fully integrated model of the tandem mirror would be developed in the future. This could entail either a 2 real-space dimension 2 velocity-space dimension (2D2V) time-dependent Fokker-Planck solver or a set of bounce-averaged 1D2V time dependent Fokker-Planck solvers coupled together with the proper boundary-conditions and source/sink terms. A 2D2V FP solver would have the added benefit of resolving many of the issues which can arise with the ambipolar potential calculation on the addition of EC heating as a more accurate method for calculating the ambipolar potential such as that proposed in Degond et al. (2010) or Kupfer et al. (1996) could be used. Such a solver would also allow the investigation of thermal barrier tandems, which could provide a large performance enhancement over the classical tandems investigated here. While it was found that the applicability of an RF ion acceleration scheme analogous to that in WHAM is limited in an end plug (see Appendix B), further work on RF heating in magnetic mirrors is warranted. This would include the development of axisymmetric mirror geometry full-wave solvers and more innovative RF schemes which reduce quasilinear particle losses. Finally, while of limited importance in a negative-ion neutral beam fueled end plug, the impact of halo neutrals generated by NBI charge exchange and edge neutral charge exchange induced loss of fast ions are anticipated to be of some importance in smaller mirrors like WHAM. To better understand these effects the modelling tools used in this study will be integrated with edge neutral modelling codes.
Acknowledgements
This research was funded by Realta Fusion and performed as part of the DOE Milestone Based Fusion Development Program DE-SC0024887. The authors would like to thank the DOE Milestone review committee for their helpful input which improved this paper. This research also used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award FES-ERCAP0026655. Realta Fusion is a fusion energy company spun out of the ARPA-E funded WHAM project at the University of Wisconsin–Madison. Realta is developing compact magnetic mirror fusion power plants as the low capital cost path to fusion power, and targeting initial fusion power plant deployment at the $100-200~{}\mathrm{MW}$ scale at industrial sites and datacenters for process heat and/or electricity.
Appendix A The effect of electron temperature variation on end plug performance
In order to investigate the effect of increased $T_{e}$ and $\phi_{p}$ on ion confinement in simple mirror end plugs, modified versions of the CQL3D-m simulations which appeared in Section 4 were used. In these simulations, CQL3D-m was run with a Maxwellian electron background species with fixed temperature and an evolving electron density set equal to the ion density. A Pastukhov potential of $\phi_{p}=4T_{e}$ was set in all simulations. This is lower than the canonical $\phi_{p}\sim 5T_{e}$ result in simple mirrors (Pastukhov, 1974) and a result of the reduced $T_{i}/T_{e}$ ratio in the tandem mirror end plug which lowers the collisionality difference between the ions and electrons. The use of this somewhat reduced $\phi_{p}$ value is supported by the calculations of $\phi_{p}$ in the POPCONs analysis in Section 3 and the literature (Logan et al., 1980; Hua & Fowler, 2004).
To facilitate efficient scoping of the impact of electron temperature and Pastukhov potential variation on performance, a fixed end plug condition was used with $B_{0}=4~{}\mathrm{T}$, $\ell=4.5~{}\mathrm{m}$, $a_{m}=0.15~{}\mathrm{m}$, and $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$. The equilibrium in these cases was left static. This is roughly consistent with an assumption that the background value of $B_{0}$ would be adjusted such that the field after diamagnetism was accounted for was $4~{}\mathrm{T}$. The velocity-space resolution in these CQL3D-m simulations could be reduced to 300 points in $u_{0}$ and 200 points in $\vartheta$ as ions and electrons did not need to be coupled, and the Pastukhov potential calculation was turned off (these calculations need a refined grid as the thermal component of both distributions and the mirror loss boundaries must be finely resolved). Similarly, the realspace $\sqrt{\psi_{n}}$ flux surface resolution could be reduced to 14 as coupling to Pleiades was not performed, and as fewer flux surfaces were used in the simulation, the number of particles in the NBI source could also be reduced to $1\times 10^{5}$ (when flux surface volumes, particularly those of the inner surfaces, are larger it becomes much easier to consistently formulate a bounce-averaged NBI source).
Figure 10: Average density $\langle n\rangle$ versus the end plug electron temperature $T_{e}$ for an endplug with $B_{m}=25~{}\mathrm{T}$, $a_{m}=0.15~{}\mathrm{m}$, $R_{m}=6.25$, $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$, $\ell=4.5~{}\mathrm{m}$, $\theta_{\mathrm{NBI}}=45^{\circ}$, and $\phi_{p}=4T_{e}$.
Values of $T_{e}$ were scanned from $25-125~{}\mathrm{keV}$ in increments of $25~{}\mathrm{keV}$ and $\phi_{p}$ was varied in a corresponding fashion to examine the effect of operating at higher temperatures and potentials than those found in the simulations described in Section 4. In addition to varying $T_{e}$, the neutral beam energy was scanned from $240-600~{}\mathrm{keV}$ in increments of $120~{}\mathrm{keV}$. The results of this scan are shown in Figure 10 and show when higher values of $T_{e}$ are present from electron heating by alphas in the central cell of the tandem, higher energy neutral beams can be much more beneficial than they are in the simple mirror case examined in Section 4. The reason for this is two-fold: high energy ions are less susceptible to losses from being dragged or scattered into the ambipolar potential, and provided $E_{\mathrm{NBI}}$ is high enough for ions to not be promptly lost to the ambipolar potential, the end plug confinement improves at high $T_{e}$ to the extent that high energy beams build up large enough plasma densities to be absorbed efficiently. It should be noted here that these CQL3D-m results draw the electron drag dominated plug assumption made in (20) to simplify the tandem mirror POPCON system into question. Luckily however, the results of (20) should only vary weakly if the plug becomes more ion-ion scattering dominated. These results also indicate that the assumption that the increased end plug electron temperature in tandem mirrors should improve confinement if slightly higher energy neutral beams are used in Section 4 is accurate, and the general takeaways from the optimization in that section still hold; a viable end plug is achievable at modest beam powers and with realistic engineering assumptions.
Appendix B Limitations of RF ion heating in end plugs
In previous high-field mirror literature, there has been some discussion of RF heating end plug ions to improve particle confinement (Forest et al., 2021; Endrizzi et al., 2023). However, it was decided to not use RF heating in this study. Over the course of this work, it became clear that an RF ion heating scheme like that employed in WHAM (Endrizzi et al., 2023), in which positive-ion, $\sim 100~{}\mathrm{keV}$, NBI ions are accelerated by $n=2-3$ harmonic ion-cyclotron heating, is not effective enough at increasing ion particle confinement, and at tandem mirror electron temperatures, positive-ion neutral beams with a beam energy of $120~{}\mathrm{keV}$ and an injection angle of $45^{\circ}$ will, according to (32), be nearly immediately lost to the ambipolar potential (this is supported by the analysis in Appendix A). In addition, particle losses from enhanced quasilinear diffusion into the loss-cone eliminate much of the gain in particle confinement that can be obtained from increasing the ion energy. While ion energy confinement and fusion rates are increased with the addition of RF heating, this is of negligible importance to overall tandem mirror fusion performance (and enhanced fusion rates in the end plug are in fact undesirable in most circumstances). In the remainder of this appendix, a representative RF heating case in at end plug conditions will be analyzed which demonstrate these challenges.
The simulation under consideration here utilizes a tritium-fueled end plug with $a_{m}=0.15~{}\mathrm{m}$, $B_{m}=25~{}\mathrm{T}$, $B_{0}=5~{}\mathrm{T}$, $\ell=4.5~{}\mathrm{m}$, $E_{\mathrm{NBI}}=120~{}\mathrm{keV}$, $P_{\mathrm{NBI}}=10~{}\mathrm{MW}$, $f_{\mathrm{RF}}=115.5~{}\mathrm{MHz}$, and $P_{\mathrm{RF}}=5~{}\mathrm{MW}$. This end plug condition is similar to the optimized end plugs identified in Section 4, but rather than using a $15~{}\mathrm{MW}$ of negative-ion NBI power, it uses $15~{}\mathrm{MW}$ combined of positive-ion NBI and $n=3$ harmonic RF power. The RF antenna in these simulations was placed slightly beyond the fast ion turning point at $z=1~{}\mathrm{m}$ calculated by CQL3D-m (this does not correspond exactly to $B/B_{0}=2$ due to the use of realistic neutral beam geometry and the effects of finite $\beta$) at the limiting flux surface $\sqrt{\psi_{n}}=1.0$, and the RF source frequency $f_{\mathrm{RF}}$ was set to correspond to a location just inside the fast ion turning point. The antenna spectrum used in GENRAY-c was consistent with the loop style of antennas which will be tested on WHAM (Endrizzi et al., 2023) and had a poloidal mode spectrum centered about $n_{\theta}=0$ with a spectral width of $n_{\theta}\pm 1$ and a uniform axial mode spectrum $n_{\phi}=0-10$ (this definition of the axial mode spectrum with $n_{\phi}$ is an artifact of the toroidal GENRAY code, but has been preserved in the text to make future replication with GENRAY-c easier). The simulations of this end plug were initially run with only NBI and no additional RF heating. CQL3D-m was iterated with Pleiades 8 times with 800 total CQL3D-m timesteps for $\sim 1~{}\mathrm{s}$ using only NBI heating until it reached steady-state. This was done in order to develop a target plasma with high enough $\beta$ to enable efficient cyclotron harmonic damping of the RF waves. When steady-state was achieved, GENRAY-c simulations were added to the CQL3D-m/Pleiades IPS workflow using self-consistent equilibrium and plasma profiles and coupled to CQL3D-m using an ion and electron quasilinear diffusion model. After the addition of RF, the simulations were run for an additional 1000 CQL3D-m time-steps, iterating Pleiades, CQL3D-m and GENRAY-c 8 times, for $\sim 1~{}\mathrm{s}$.
Figure 11: On the left ion density (blue) and temperature (red) profiles versus the normalized square root poloidal flux $\sqrt{\psi_{n}}$ for a simulation with RF (solid) and without RF (dashed), and on the right the deposited RF power density profiles versus $\sqrt{\psi_{n}}$.
The radial plasma $\langle E_{i}\rangle$ and $n$ profiles as well as the RF power deposition profiles from the simulations described above are shown in Figure 11. The addition of RF heating did provide a slight increase in the overall particle confinement and particle density, but the overall plug performance remained well below that which could be obtained using $15~{}\mathrm{MW}$ of negative-ion NBI (reference the plot shown in Figure 9 to see a comparable negative-ion NBI case). The majority of the improvement in particle confinement appears to have come from the $T_{e}$ increase due to the $\sim 1~{}\mathrm{MW}$ of electron directed RF heating by Landau damping. Ion directed heating’s effect on plasma performance was a mixed. There was a significant amount of inverse cyclotron harmonic damping that diffused particles downwards in velocity-space into the loss cone. Furthermore, quasilinear diffusion cannot increase the density anywhere in velocity-space to a larger value than it is at the neutral beam source location. Thus, particles at the neutral beam source energy continued to dominate the overall velocity-space pitch-angle scattering rate. Energy confinement increased, with the plasma reaching high average ion energies due to the formation of an RF driven energetic ion tail and quasilinear diffusion driven pumpout of lower energy ions, however, this is actually undesirable in an end plug. Just increasing average end plug particle energy without improving end plug particle confinement will not improve the tandem mirror central cell’s performance and will only serve to drive up $\beta$ putting the end plug closer to pressure driven instability limits. One potential solution would be to use higher harmonics which bias the magnitude of the quasilinear diffusion operator upwards in velocity space due to the quasilinear diffusion operator’s Bessel function dependence $Q\propto\mathrm{J}_{n}^{2}(k_{\perp}v_{\perp}/n\Omega_{i})$. However, harmonics greater than $n=3$ were found to be weakly damped in mirror end plug plasmas. Weak damping is undesirable in reactor scenarios as it can lead to poor antenna coupling and large in vessel fields that generate impurity sources. To increase damping rates, lower harmonics, $n=2,3$, had to be used. At lower harmonics, after the fast ion tail built up, damping became relatively strong, as shown in Figure 12, and the rays in the simulation damped after two passes through the resonance.
Figure 12: Plots of rays in black versus $r$ and $z$. Resonances are indicated with the rainbow colored labeled contours on the plot. Poloidal flux, $\psi$, surfaces are shown in blue.
As a result of simulations like the representative case shown here, RF ion heating was dropped from this study. The WHAM RF ion heating scheme was determined to not be scalable to tandem mirror fusion reactor end plugs. Positive-ion NBI cannot effectively fuel a tandem mirror end plug at relevant values of $T_{e}$ and $\phi_{p}$ and RF quasilinear diffusion at $n=2,3$ harmonics with this scheme was not effective enough at increasing particle confinement times. Finally, in cases of both weak and the strong damping there could potentially be validity issues with the raytracing approximation here. In cases of weak damping, interference and diffraction may impact the wave power deposition, and in cases of strong damping, the distance between the antenna and the wave resonance are comparable to the wavelength of the wave. Because of this, full-wave simulations may be needed to adequately model RF in magnetic mirrors due to the breakdown of the raytracing approximation. Future work may revisit RF ion heating in conjunction with negative ion neutral beams, modified antenna launch spectra which optimize the shape of quasilinear diffusion operator and the use of a full wave code such as AORSA (Jaeger et al., 2002).